# Stochastic dominance for super heavy-tailed random variables

### v1

@@ 3.1 Properties of super heavy-tailed distributions
$G\left(\frac{xy}{x+y}\right)\geq G(x)G(y)~{}~{}~{}\mbox{for all $x,y>0$.}$ (4)
@@
Hence, (4) holds for $n=2$. Next, assume that (4) holds for $n=k-1$ where $k>3$ is an integer. Let $a=\sum_{i=1}^{k-1}\theta_{i}F_{i}(x)$, $b=\sum_{i=1}^{k-1}\theta_{i}F_{i}(y)$, $c=a/(F_{n}(x)(1-\theta_{n}))$, and $d=b/(F_{n}(y)(1-\theta_{n}))$. For $n=k$,
@@
As $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{k}$, $F_{k}\leq\sum_{i=1}^{k-1}\theta_{i}/(1-\theta_{n})F_{i}$. Thus $c,d\geq 1$ and $cd+1-c-d\geq 0$. The proof is complete by induction. ∎
Remark 2.
Let $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ be some function. It is well-known that if $f(x)/x$ is decreasing, $f$ is sub-additive (Theorem 7.2.4 of Hille and Phillips, (1996)). Another sufficient condition for $f$ being sub-additive is that $f$ is concave and $\lim_{x\downarrow 0}f(x)\geq 0$ (Theorem 7.2.5 of Hille and Phillips, (1996)). Hence, if $h_{F}(x)/x$ is decreasing or $h_{F}$ is concave, $F$ is super heavy-tailed.
3.2 Examples of super heavy-tailed distributions
In this section, we demonstrate that many well-known infinite-mean distributions belong to the super heavy-tailed class.
Example 1 (Pareto distribution).
@@ 3.2 Examples of super heavy-tailed distributions
$\mathbb{P}(X\leq x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
If $\alpha\leq 1$, the Pareto distribution is super heavy-tailed. This can be seen by noting that any Pareto$(\alpha)$ random variable $X$ can be written as $X=f(Z)$ where $Z\sim{\rm Pareto}(1)$ and $f(x)=(x+1)^{1/\alpha}-1$ for $x\geq 0$. It is easy to show that Pareto$(1)$ is super heavy-tailed. As $f$ is increasing and convex if $\alpha\leq 1$, the result follows from Proposition 2 (iv).
Example 2 (Fréchet distribution).
For $\alpha>0$, the Fréchet distribution is defined as
$\mathbb{P}(X\leq x)=\exp(-x^{-\alpha}),~{}~{}x>0.$
If $\alpha\leq 1$, the Fréchet distribution is super heavy-tailed. To see this, let $F$ denote the distribution function of $X$. We have $h_{F}(x)=x^{\alpha}$ for $x>0$. Then, if $\alpha\leq 1$, for $x,y>0$,
Thus $h_{F}$ is sub-additive and the Fréchet distribution with $\alpha\leq 1$ is super heavy-tailed.
Example 3 (Generalized Pareto distribution).
@@
$\mathbb{P}(X\leq x)=\begin{cases}1-\left(1+\xi\frac{x}{\beta}\right)^{-1/\xi}&x\in D(\xi,\beta)\\ e^{-x/\beta}&x\geq 0,\end{cases}$
where $D(\xi,\beta)=[0,\infty)$ if $\xi>0$ and $D(\xi,\beta)=[0,-\beta/\xi)$ if $\xi<0$. By the Pickands-Balkema-de Haan Theorem (Balkema and de Haan,, 1974; Pickands,, 1975), the generalized Pareto distributions are the only possible non-degenerate limiting distributions of the excess of random variables beyond a high threshold. If $\xi\geq 1$, then the generalized Pareto distribution is super heavy-tailed. This is by Proposition 2 (iv); that is, the generalized Pareto random variables with $\xi\geq 1$ can be obtained from location-scale transforms of Pareto$(1/\xi)$ random variables.
Example 4 (Super-Pareto distribution).
Introduced by Chen et al. (2024a), a super-Pareto distribution is the distribution of a random variable $f(X+1)$ where $X\sim\rm{Pareto}(1)$ and $f$ is some non-decreasing, convex, and non-constant function. By Proposition 2 (iv), a super-Pareto distribution (after some location-scale transform) is super heavy-tailed.
Example 5 (Burr distribution).
@@
$\mathbb{P}(X\leq x)=1-\left(\frac{1}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0.$ (5)
Let $Y\sim\rm{Pareto}(\alpha)$. Then $Y^{1/\tau}$ follows a Burr distribution. If $\alpha,\tau\leq 1$, the Burr distribution is super-Pareto and hence super heavy-tailed. Special cases of Burr distributions are the paralogistic ($\alpha=\tau$) and the log-logistic ($\alpha=1$) distributions; see Kleiber and Kotz, (2003) and Klugman et al. (2012).
Example 6 (Inverse Burr distribution).
Suppose that $Y$ follows the Burr distribution (5). Then $X=1/Y$ follows the inverse Burr distribution
$\mathbb{P}(X\leq x)=\left(\frac{x^{\tau}}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0,$
where $\alpha,\tau>0$. Let $F$ denote the inverse Burr distribution function. If $\tau\leq 1$, it is easy to check that the second derivative of $h_{F}$ is always negative, and thus $h_{F}$ is sub-additive (see Remark 2). Hence $X$ is super heavy-tailed if $\tau\leq 1$. Note that the super heavy-tailed property may not always be preserved under the inverse transformation of super heavy-tailed random variables. For instance, if $Z$ follows a Fréchet distribution without finite mean, then $1/Z$ follows a Weibull distribution whose mean is always finite and it cannot be super heavy-tailed.
Example 7 (Log-Pareto distribution).
@@
$\mathbb{P}(X\leq x)=1-\frac{1}{(\log(x+1)+1)^{\alpha}},~{}~{}x>0.$
If $\alpha\in(0,1]$, by Proposition 2 (iv), $X$ is super heavy-tailed.
Example 8 (Stoppa distribution).
@@
$\mathbb{P}(X\leq x)=\left(1-\frac{1}{(x+1)^{\alpha}}\right)^{\beta},~{}~{}x>0.$

### v3

@@ Appendix B Proof of Proposition 3
$G\left(\frac{xy}{x+y}\right)\geq G(x)G(y)~{}~{}~{}\mbox{for all $x,y>0$.}$ (A.2)
@@
Hence, (A.2) holds for $n=2$. Next, assume that (A.2) holds for $n=k-1$ where $k>3$ is an integer. Let $a=\sum_{i=1}^{k-1}\theta_{i}F_{i}(x)$, $b=\sum_{i=1}^{k-1}\theta_{i}F_{i}(y)$, $c=a/(F_{n}(x)(1-\theta_{n}))$, and $d=b/(F_{n}(y)(1-\theta_{n}))$. For $n=k$,
@@
As $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{k}$, $F_{k}\leq\sum_{i=1}^{k-1}\theta_{i}/(1-\theta_{n})F_{i}$. Thus $c,d\geq 1$ and $cd+1-c-d\geq 0$. The proof is complete by induction.
Appendix C Examples of distributions in the class $\mathcal{H}$
In this section, we demonstrate that many well-known infinite-mean distributions are in $\mathcal{H}$.
Example A.1 (Pareto distribution).
@@ Appendix C Examples of distributions in the class $\mathcal{H}$
$F(x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
If $\alpha=1$, for $x,y>0$,
$h_{F}(x+y)-h_{F}(x)-h_{F}(y)=\log(x+y+1)-\log(x+1)-\log(y+1)\leq 0$
Thus, Pareto$(1)\in\mathcal{H}_{s}$. Then we note that any Pareto$(\alpha)$ random variable $X$ can be written as $X=f(Z)$ where $Z\sim{\rm Pareto}(1)$ and $f(x)=(x+1)^{1/\alpha}-1$ for $x\geq 0$. By Proposition 2 (iv), as $f$ is increasing and convex for $\alpha\leq 1$, Pareto$(\alpha)\in\mathcal{H}_{s}$ if $\alpha\leq 1$.
Example A.2 (Generalized Pareto distribution).
@@
$F(x)=\begin{cases}1-\left(1+\xi\frac{x}{\beta}\right)^{-1/\xi},&\mbox{~{}~{}if$\xi\neq 0$},\\ e^{-x/\beta},&\mbox{~{}~{}if $\xi=0$},\end{cases}$
where $x\in[0,\infty)$ if $\xi\geq 0$ and $x\in[0,-\beta/\xi)$ if $\xi<0$. By the Pickands-Balkema-de Haan Theorem (Balkema and de Haan,, 1974; Pickands,, 1975), the generalized Pareto distributions are the only possible non-degenerate limiting distributions of the excess of random variables beyond a high threshold. If $\xi\geq 1$, $F\in\mathcal{H}$. This is by Proposition 2 (iv); that is, the generalized Pareto random variables with $\xi\geq 1$ can be obtained from location-scale transforms of Pareto$(1/\xi)$ random variables.
Example A.3 (Burr distribution).
@@
$F(x)=1-\left(\frac{1}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0.$ (A.3)
Let $Y\sim\rm{Pareto}(\alpha)$. Then $Y^{1/\tau}$ follows a Burr distribution. If $\alpha,\tau\leq 1$, the Burr distribution is super-Pareto and hence $F\in\mathcal{H}$. Special cases of Burr distributions are the paralogistic ($\alpha=\tau$) and the log-logistic ($\alpha=1$) distributions; see Kleiber and Kotz, (2003) and Klugman et al. (2012).
Example A.4 (Inverse Burr distribution).
Suppose that $Y$ follows the Burr distribution (A.3). Then $X=1/Y$ follows the inverse Burr distribution
$F(x)=\left(\frac{x^{\tau}}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0,$
where $\alpha,\tau>0$. If $\tau\leq 1$, it is easy to check that the second derivative of $h_{F}$ is always negative, and thus $h_{F}$ is subadditive. Hence $F\in\mathcal{H}$ if $\tau\leq 1$. Note that the property of $\mathcal{H}$ may not always be preserved under the inverse transformation. For instance, if $Z$ follows a Fréchet distribution without finite mean, then $1/Z$ follows a Weibull distribution whose mean is always finite.
Example A.5 (Log-Pareto distribution).
@@
$F(x)=1-\frac{1}{(\log(x+1)+1)^{\alpha}},~{}~{}x>0.$
If $\alpha\in(0,1]$, by Proposition 2 (iv), $F\in\mathcal{H}$.
Example A.6 (Stoppa distribution).
@@
$F(x)=\left(1-\frac{1}{(x+1)^{\alpha}}\right)^{\beta},~{}~{}x>0.$